OpenStudy (anonymous):

Solve y''+y=cot(x). I know I asked a similar, I had the signs incorrect, my bad. Ok from here I have the characteristic equation is r^2+1=0 which gives r=i,-i. this give y1=cosx and y2=sinx.....right? Where do I go from here?

OpenStudy (anonymous):

Like last time, I'm pretty sure you'll need to use variation of parameters. I haven't come across a solution to an equation like this that did not involve VoP, other than (maybe) power series.

OpenStudy (anonymous):

Apparently we DID learn it but he never ever gave it a name in class. I had to find it in the book and it's basically like what he did in examples. It was right after undetermined coefficients. I'm sorry I didn't realize that yesterday!

OpenStudy (anonymous):

That's okay! First thing you do is check the fundamental solutions \(\sin x\) and \(\cos x\) are linearly independent. Do you know how to do that with the Wronskian?

OpenStudy (anonymous):

[cosx sinx -sinx cosx] and then take the determinant?

OpenStudy (anonymous):

and the determinant=1 which isn't 0 so they are linearly independent....?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Now you use the formulas for the new-parameter solution, \[u_1=-\int\frac{y_2g(t)}{W(y_1,y_2)}~dt~~~~\text{and}~~~~u_2=\int\frac{y_2g(t)}{W(y_1,y_2)}~dt\] where \(y_1=\cos x\), \(y_2=\sin x\), \(W(y_1,y_2)\) is the Wronskian, and \(g(t)=\cot x\). The final solution will be \[y=u_1y_1+u_2y_2\]

OpenStudy (anonymous):

oh man....the answer's gonna be ugly isn't it?

OpenStudy (anonymous):

Surprisingly, no! Well, maybe a little, but one of those integrals will simplify to something really nice.

OpenStudy (anonymous):

should that be y1 in the formula for u1?

OpenStudy (anonymous):

Nope, \(u_1\) is right, \(u_2\) is wrong: \[u_1=-\int\frac{y_2g(t)}{W(y_1,y_2)}~dt~~~~\text{and}~~~~u_2=\int\frac{\color{red}{y_1}g(t)}{W(y_1,y_2)}~dt\] Thanks for catching that.

OpenStudy (anonymous):

so.....sorry I'm just making sure I have all steps correct.....u1 should become -sin(x)...and u2= cos(x) +log(sin(x/2))-log(cos(x/2))

OpenStudy (anonymous):

and the final, general solution will be c1cosx +c2sinx + (u1y1 +u2y2)?

OpenStudy (anonymous):

\(u_1\) is correct. I don't have any paper at hand so I'll check your work with the second integral: \[\begin{align*}u_2&=\int\cos x\cot x~dx\\&=\cot x\sin x+\int\csc^2x\sin x~dx&\text{integrating by parts}\\ &=\cos x+\int\csc x~dx\\ &=\cos x-\ln\left|\csc x+\cot x\right| \end{align*}\] With some identities, you'd get \[u_2=\cos x+\ln\left|\sin\frac{x}{2}\right|-\ln\left|\cos\frac{x}{2}\right|\] So yes, you're right. And yes, the general solution would be as you wrote it.

OpenStudy (anonymous):

AWESOME!!!! thank you so so so much!

OpenStudy (anonymous):

You're welcome!